OpenStudy (anonymous):

What is 14.2% expressed as a decimal?

OpenStudy (anonymous):

.142

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (anonymous):

@zepdrix

OpenStudy (anonymous):

why

OpenStudy (alexandervonhumboldt2):

\[\frac{ 14.2 }{ 100 }=\frac{ 142/10 }{ 100 }=0.142\]

OpenStudy (alexandervonhumboldt2):

sorry, you was correct

OpenStudy (anonymous):

k

OpenStudy (anonymous):

medal 4 u